OpenStudy (anonymous):

Solve this Linear DE: y'+3x^2y=x^2 need a check at the end

OpenStudy (anonymous):

\[e^{\int{}{}p(x)dx}=e^{\int{}{}3x^2}=e^{x^3}\]

OpenStudy (anonymous):

\[\frac{d}{dx}e^{x^3}y=e^{x^3}x^2\]

OpenStudy (anonymous):

\[e^{x^3}y=\int{}{}e^{x^3}x^2dx\]

OpenStudy (anonymous):

u=x^3 du=3x^2dx du/3=x^2dx

OpenStudy (anonymous):

\[e^{x^3}y=\frac{1}{3}\int{}{}e^udu\]

OpenStudy (anonymous):

\[e^{x^3}y=\frac{1}{3}e^{u}+\frac{1}{3}c=\frac{e^{x^3}}{3}+\frac{c}{3}\]

OpenStudy (anonymous):

c/3 = c

OpenStudy (anonymous):

\[y=\frac{1}{3}+ce^{-x^3}\]

OpenStudy (anonymous):

I chcked it was correct

OpenStudy (anonymous):

looks right to me